OpenStudy (anonymous):

What is the sum of the arithmetic sequence 140, 133, 126…, if there are 33 terms? 689 896 924 782

OpenStudy (slaaibak):

Tn = a + (n-1)d T1= 140 = a d = 133 - 140 = -7 T33 = 140 + 32*-7 = -84 Sum = n/2(first + last) = 16.5(140 - 84) = 824

OpenStudy (slaaibak):

Sorry, mistyped that, it's 924

OpenStudy (anonymous):

Thank you